ganeshie8 (ganeshie8):

unit's digit of \(\large 3^{3^{3^{3^{3^{3^{3}}}}}} \)

OpenStudy (anonymous):

awww to many 3's

OpenStudy (anonymous):

chooo many 3's :3

ganeshie8 (ganeshie8):

7 threes only :)

OpenStudy (anonymous):

well hmm let me think of a battern instead of calculation

OpenStudy (anonymous):

3, 7, 1, 3, 7, 1, 3, 7, 1 this will be the last digits for every corresponding powers..

OpenStudy (anonymous):

I think 3 will the last digit...

OpenStudy (anonymous):

3=3 mod 10 3^2= 9 mod 10 3^3=7 mod 10 3^4=1 mod 10 3^5=3 mod 10 ohk fairr enough :3 4k+q

ganeshie8 (ganeshie8):

3, 9, 7, 1...

OpenStudy (anonymous):

9 ??

OpenStudy (anonymous):

so next step seeying 3^3^3^3^3^3 mod 4

OpenStudy (anonymous):

Wait, let me think more, I am wrong..

OpenStudy (anonymous):

hm 3=-1 mod 4 thus 3^(odd ) =-1 mod 4 therfore checking for 4k+3 which gives 1 yeahhhh

OpenStudy (anonymous):

:3

ganeshie8 (ganeshie8):

wolfram says 7 though http://www.wolframalpha.com/input/?i=3%5E3%5E3%5E3%5E3%5E3%5E3%5E3

OpenStudy (anonymous):

3 and 7 will repeat..

OpenStudy (anonymous):

xD

ganeshie8 (ganeshie8):

3,9,7,1 so it repeats every fourth power

ganeshie8 (ganeshie8):

how did u conclude "1" @BSwan

ganeshie8 (ganeshie8):

``` 3=-1 mod 4 thus 3^(odd ) =-1 mod 4 therfore checking for 4k+3 ``` makes sense so far

ganeshie8 (ganeshie8):

you get 3^(4k+3) finally, right ?

OpenStudy (anonymous):

wait ahahahaha yeah yeah which gives 7 :3

OpenStudy (anonymous):

xD i only counted 0,1,2,3 insted of 1,2,3,0 xD from battern

OpenStudy (anonymous):

i love 7 anyway its hard number

ganeshie8 (ganeshie8):

Ahh that makes sense :) so basically you have worked it something like below : 3^3^3^3^3^3^3 = 3^(4k+3) = 3^3*(3^4)^k = 7(1)^k = 7 nice :)

OpenStudy (anonymous):

awww now i understand if its -1 why it keeps being -1 :o all odd powers ! dint relize that on noon , im feel im much better now lol

OpenStudy (anonymous):

hehehehe

OpenStudy (anonymous):

it does water in eyes :D

ganeshie8 (ganeshie8):

i would have added more 3's proabably haha!

myininaya (myininaya):

I'm seeing a 7,3,7,3,... type of pattern

OpenStudy (anonymous):

why there is water in eyes anyway ? that remindes me with a novel memories of geisha , did u read it ?

OpenStudy (anonymous):

There are 8 3s in wolfram... But answer is yet the same..

ganeshie8 (ganeshie8):

@myininaya we need to work it from top/exponent right, pattern might be tough to predict for the last digit i thinik

OpenStudy (anonymous):

No I have not @BSwan

OpenStudy (anonymous):

why the answer is same cuz 3=-1 mod 4 thus 3^(any odd power )=-1 mod 4

ganeshie8 (ganeshie8):

3^3^3^3^3^3^3 = 3^(3^3^3^3^3) its not same as (3^3)^3^3^3^3 or something... not so sure, need to check...

myininaya (myininaya):

oh you know what i was thinking about the first digit my bad

ganeshie8 (ganeshie8):

Oh! never thought of left most digit before xD last digit is easy to access using mod 10... but left most digit might be tough to work

OpenStudy (anonymous):

left most digit ?

myininaya (myininaya):

wait

OpenStudy (anonymous):

hmm well find a battern of (3*10)^ something hehe cool lets think of it

myininaya (myininaya):

isn't the first digit the ones digit and the last digit the left-most digit

OpenStudy (anonymous):

I am no good at MOD.. we were not taught this topic in mathematics anywhere in Schooling.. :)

myininaya (myininaya):

i'm confused about those terms first and last

OpenStudy (anonymous):

Last digit is one's digit..

ganeshie8 (ganeshie8):

Oh sorry, i should have used "unit's digit"

OpenStudy (anonymous):

hmm shure ?

ganeshie8 (ganeshie8):

units tens hundred thousands... yes sure, this is the correct terminologhy :)

OpenStudy (anonymous):

lavosh >.<

myininaya (myininaya):

well if it really was trying to figure out the units digit then it would be easy right i think \[3^3=..7\] \[3^{3^3}=(..7)^3=\] well since 7*7*7=49*7=...3 then we know \[3^{3^3}=(..7)^3=...3\] \[3^{3^{3^3}}=(...3)^3 \] 3*3*3=9*3=27 so we know \[3^{3^{3^3}}=(...3)^3=...7 \] this is how i got my 7,3,7,3,7,3... pattern

myininaya (myininaya):

I hope you guys know those ... just mean blah blah

OpenStudy (anonymous):

hmm ok -.- makes me feel we did nothing up there

OpenStudy (anonymous):

i thought ur gonna comment something fantacy like last digit >.< lavosh girl

myininaya (myininaya):

No no I wasn't trying to imply you guys were doing nothing

ganeshie8 (ganeshie8):

Now I see :) but you seem to be starting from the bottom term first ? \[ \large \color{Red}{3}^{\color{red}{3}^{3^{3^{3^{3^{3}}}}}} \]

myininaya (myininaya):

I was

myininaya (myininaya):

or I did whatever

ganeshie8 (ganeshie8):

i see haha! i think it would be bit more hard if we had to work it from the top first

myininaya (myininaya):

I think these type of problems aren't normally looked at the way I approached but with the whole mod thing

ganeshie8 (ganeshie8):

because 3^(3^3) gives 3^27 ****Bswan method**** (3^3)^3 gives 27^3 ****your method****

ganeshie8 (ganeshie8):

your method is easy to work because we can take mod 10 readily for the bottom 27

ganeshie8 (ganeshie8):

Bswan's method is bit more challenging as we don't know what happens in the exponent

OpenStudy (anonymous):

challenging ? hmm

OpenStudy (anonymous):

well , its not you who can juge :P

ganeshie8 (ganeshie8):

i mean, compared to (27)^3, finding the unit's digit of 3^27 is challenging

OpenStudy (anonymous):

yeah right using simple theory is much chalenging than calculating 27^3 lol redicules jugment

OpenStudy (anonymous):

what if u have already base of 27^27^27^27 ?

ganeshie8 (ganeshie8):

what i meant is, working units digit of some number like : (3408504385094385)^3 is easy compared to 3^3^3^3^3^3

ganeshie8 (ganeshie8):

since it is a single exponent, you can work it in a snap

OpenStudy (anonymous):

ur opinion , its upto what you know and what u could memorize to solve something , i wont juge in anyway both looks cool to me , but i wont say something is better cuz bla blah blah

ganeshie8 (ganeshie8):

however we cannot start from base for working the unit's digit of 3^(3^(3^(3^(3^(3^3)))))

myininaya (myininaya):

I don't think he was saying one way was better than the other

myininaya (myininaya):

just because something is found to be more challenging doesn't mean it is a worse way it means to me anyways that it is a way worth exploring

OpenStudy (anonymous):

see ,it depands on u :P

ganeshie8 (ganeshie8):

exactly ! @BSwan i was saying the problems are different, so they both require different methods. your method works for : 3^(3^(3^(3^(3^(3^3))))) myininaya's method works for : ((((((3^3)^3)^3)^3)^3)^3

ganeshie8 (ganeshie8):

they're two different problems, and two different methods. your method and the expression are challenging myininaya's method and the expression are simpler

OpenStudy (anonymous):